OpenStudy (anonymous):

7x^2 + xy + 7y^2 = 15, (1, 1) (ellipse) find the tangent line at the given point (1,1) by implicit differentiation

OpenStudy (aum):

Can you do the implicit differentiation part?

OpenStudy (phi):

I would do \[ \frac{d}{dx}(7x^2 + xy + 7y^2 = 15) \\ 7\frac{d}{dx} x^2 + \frac{d}{dx}(x y) + 7 \frac{d}{dx} y^2 = \frac{d}{dx} 15\] can you do *any* of those terms?

OpenStudy (anonymous):

y=-x+2

OpenStudy (aum):

correct.